2. Suppose that a program does read operations on the following memory addresses (e.g., with “lb” or “lw” instructions): 248, 1312. Give the number of the memory block that each of these addresses belongs to, for each of the following memory block sizes.

(a) block size of one word (4 bytes)

(b) block size of eight words (32 bytes)

3. Give the position (or set) in the cache that would be checked on each of the read operations of question 2, for each of the following caches.

(a) direct-mapped cache with total capacity of 32 one-word blocks

(b) direct-mapped cache with total capacity of 4 eight-word blocks

(c) 4-way set-associative cache with total capacity of 32 one-word blocks

(d) 2-way set-associative cache with total capacity of 4 eight-word blocks
Answer #1

ANSWER:

GIVEN THAT:

In the byte addresable memory, 'lb' or 'lw' instructions fetches one byte from memory

a) block size of one word(4 Bytes)

i.e 4 Bytes are present in one memory block, So, for given addresses

i) 248 belongs to 248/4 = 62 block

ii) 1312 belongs to 1312/4=328 block

b) block size of eight word (32 Bytes)

So,

i) 248 belongs to 248/32=ceil(7.75) = 8 block

ii) 1312 belongs to 1312/32=41 block

Note that these blocks are indexed from 0.